OpenStudy (anonymous):

Which equations represent the line shown in the graph? Choose all answers that are correct.

OpenStudy (anonymous):

I think it's 1,3,and 4. I didn't solve for 2 though

OpenStudy (danjs):

recall the slope-intercept y - y1 = m(x-x1) where (y1,x1) is a point on the line m - is the slope of the line

OpenStudy (danjs):

can you calculate the slope from the 2 given points? (3,2) and (-2,-1)

OpenStudy (ahsome):

What you would need to do is to find the equation given the two points. Firstly, lets find the slope of the equation \[m=\frac{y_1-y_2}{x_1-x_2}\] Where \((x_1,y_1)\) is one co-ordinate, and \((x_2,y_2)\) Is another co-ordinate. Here, lets make: \((x_1,y_1)\)=(3,2) \((x_2,y_2)\)=(-2,-1) Lets sub these values into the equation: \[m=\frac{y_1-y_2}{x_1-x_2}\] \[m=\frac{2-(-1)}{3-(-2)}\] \[m=\frac{3}{5}\] Now, we know that the basic equation is \[y-y_1=m(x-x_1)\] So now, lets sub in the values we know \[y-y_1=m(x-x_1)\]\[y-3=\frac{3}{5}(x-2)\] That is the same as the first one you showed, so thats the answer

OpenStudy (ahsome):

That should help, @zlamar11

OpenStudy (anonymous):

thanks man big help

OpenStudy (ahsome):

No probs man :)